Suppose that you and a friend are playing cards and you decide to make a friendly wager. The bet is that you will draw two cards without replacement from a standard deck. If both cards are diamonds, your friend will pay you $801. Otherwise, you have to pay your friend $48. Step 1 of 2 : What is the expected value of your bet? Round your answer to two decimal places. Losses must be expressed as negative values.
Answer #1

let X be the amount of net gain to you

P(X=801)=P(both cards of diamond) =P(first diamond)*P(2nd diamond |first diamond)

=(13/52)*(12/51)

P(X=-48) =P*(not both are diamond) =(1-(13/52)*(12/51))

therefore expected value of bet =E(X) =xP(X) =801*((13/52)*(12/51))+(-48)*(1-(13/52)*(12/51))

= 1.94